# WebVoyager [Uncaptioned image]: Building an End-to-End Web Agent with Large Multimodal Models

### v3

@@ 1 Introduction
Figure 1: The overall workflow of WebVoyager. WebVoyager takes web tasks assigned by a human and automatically browses the web online. At each step, WebVoyager selects actions based on screenshots and text (the ‘type’ of the web element and its contents). Once the task is completed, the answers will be returned to the user. For example, a user queries: "Find the cost of a 2-year protection for PS4 on Amazon." Agent interacts with Amazon online, locates the PS4, identifies the 2-year protection price, and returns "\$30.99" to the user.
@@
We conduct evaluations on a newly-collected dataset, which is semi-automatically generated using a self-instruct (Wang et al., 2022) method, comprising 643 web tasks from 15 commonly accessed websites. We also evaluate WebVoyager on 90 web-related tasks of level 1 and level 2 from the GAIA (Mialon et al., 2023), and 50 interactive open-web tasks from SeeAct Zheng et al. (2024). We compare our WebVoyager with 1) GPT-4 (All Tools) [^2] GPT-4 (All Tools) is an integrated tool-based agent released by OpenAI in Oct. 2023. See https://chat.openai.com/, and 2) WebVoyager in a text-only setting which employs the textual accessibility tree proposed in WebArena (Zhou et al., 2023) to describe web pages. The results show that WebVoyager achieves a Task Success Rate of 59.1% on our new benchmark, significantly outperforming GPT-4 (All Tools) with a rate of 30.8% and the text-only setting with a rate of 40.1%, demonstrating the effectiveness of our method. Our research demonstrates the effectiveness of the WebVoyager method for web tasks, offering insights into the development of more intelligent and efficient web automation solutions.
@@ 2 Related Work
Along with these new benchmarks, numerous efforts have been made to build autonomous web agents. WebGPT (Nakano et al., 2021) constructs a text-based web browsing environment and fine-tunes GPT-3 as a web agent. WebAgent (Gur et al., 2023) pretrains a T5 model to extract HTML snippets and leverages Flan-U-PaLM (Chowdhery et al., 2023) to generate Python code to interact with the environment. Besides fine-tuning, another line of work tries to build web agents by prompting LLMs Yao et al. (2022b); Shinn et al. (2023); Ma et al. (2023). Multimodal web agents that integrate visual signals have also been explored, WebGUM (Furuta et al., 2023) combines T5 (Raffel et al., 2020) with a Vision Transformer (ViT) to navigate using both screenshots and HTML text. PIX2ACT (Shaw et al., 2023) instead solely relies on web screenshots as inputs to predict agent actions. Unlike previous works that only consider a single modality or simplified web environments, we build a multimodal agent that can complete tasks on real-world websites in this work. Concurrently with our work, SeeAct (Zheng et al., 2024) also leverages Large Multimodal Models (LMMs) for integrated visual understanding and actions on websites. However, the best SeeAct agent still relies on a finetuned cross-encoder model to select candidate elements for interaction. In contrast, WebVoyager do not require any additional modules.
@@ 3.3 Observation Space
Similar to how humans browse the web, our agent also takes the visual information from the web (screenshots) as the primary source of input. Using screenshots allows us to avoid the burden of processing HTML DOM tree or accessibility tree to portray the overall structure of webpages, which can lead to overly verbose texts and impact the decision-making process of the agent. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we overlay bounding boxes of the interactive elements on the websites to better guide the agent’s action prediction. Unlike Yang et al. (2023a), we do not need any object detection module (Zou et al., 2023). Instead, we utilize GPT-4-ACT [^5] https://github.com/ddupont808/GPT-4V-Act, a Javascript tool to extracts the interactive elements based on web element types and then overlays bounding boxes with numerical labels on the respective regions of the elements.
@@ 3.4 Action Space
We define the action space of our agent similar to how human browse the web. To this end, we implement the most commonly used mouse and keyboard actions, sufficient for the agent to browse various web pages and locate the content required for the task. With the help of numerical labels in screenshots, we enable the agent to respond with a concise Action Format. This approach precisely locates the elements requiring interaction and executes the corresponding actions. The usage of actions is as follows (more details in Appendix C): 1) Click. This action involves clicking on an element within a webpage, typically a link or a button. 2) Input. This composite action involves selecting a text box, deleting any existing content within it, and then inputting new content. 3) Scroll. Scrolling is a common operation for browsing webpages, usually involving the vertical movement of the entire page. 4) Wait. Action execution requires time, and this action is often used to wait for web pages to load. 5) Back. This action is used to return to the previous page. 6) Jump to Search Engine. There are often situations where agents get stuck at a certain website without finding an answer. This action enables the agent to jump to a search engine and start anew. 7) Answer. Once all questions in the task are resolved, this action concludes the iteration and provides an answer in line with the task requirements.
@@ 5 Experiment
We employ GPT-4 Turbo with vision (gpt-4-vision-preview) as the backbone model of our agent, which showcases strong semantic and visual understanding capabilities equivalent to GPT-4V (OpenAI, 2023). We include the GPT-4 (All Tools) for baselines, which integrates vision, web browsing, code analysis, and various plugins in one model. Additionally, we consider a text-only baseline where the agent only receives the websites’ accessibility tree as input to prediction actions. We used a fixed browser window size of 1024 * 768 pixels for our environment, ensuring a consistent size for the screenshots in our observations. We set the temperature to 1 during generation and allow the agent to explore at most 15 steps.
@@ 5.1 Evaluation Methods
We adopt human evaluation as our main evaluation metric since most of the questions in our benchmark have open-ended answers. In particular, we provide the human evaluators the complete trajectories of the agent’s interaction with the web (all screenshots and all actions), and ask them to provide a binary judgement of whether the agent successfully completed the task. For a subset of 300 tasks, we invite three annotators to judge each trajectory to understand the agreement among human annotators.
Even though human evaluations are accurate, they are often not scalable. Hence, we want to see if leveraging a LMM for automatic evaluation is feasible. To this end, we propose to use GPT-4V as an auto-evaluator that emulates the behavior of human evaluators to evaluate the navigation trajectories of WebVoyager. In particular, we provide the task, the responses from WebVoyager, and the last $k$ screenshots to the evaluator and ask it to judge whether the agent has completed the task, where $k$ is a hyper-parameter. The prompt of the GPT-4V evaluator is shown in Appendix B.
@@ 5.2 Result
We report the Agreement (the ratio of overlap) and Kappa ($\kappa$; Cohen 1960) between consolidated human labels666the Fleiss’s Kappa (Fleiss, 1971) of human annotators before any discussion is 0.7, which is substantial agreement. and GPT-4V’s judgments on the subset of 300 tasks in Table 2. Here, $k$ denotes the number of screenshots provided to GPT-4V, with “full” implying the full trajectory. GPT-4V’s agreement with human annotators gradually improves as it receives more information, and its final Kappa score also reaches 0.7, which is on par with the agreement among human annotators. The consistency between GPT-4V and humans suggests that GPT-4V is a promising automatic evaluator for multi-modal web agents.
@@ 5.3 Discussions
Both text and vision are necessary for generalist web agents. As discussed earlier, WebVoyager struggles with text-heavy websites. On the other hand, we observe that text-only agent demonstrates significantly poorer performance on websites with complex visual elements, such as Booking and Flights, which require interactions with calendars and other intricate components. In these scenarios, the textual input such as accessibility tree becomes highly complex and verbose, making it far less intuitive than using screenshots. Hence it’s necessary to incorporate both modalities of inputs when building the general purpose agents.
@@ 6 Conclusion
We introduce WebVoyager, an innovative web agent powered by large multimodal models (LMM) that can complete real-world web tasks end-to-end by interacting with websites. We have shown through evaluations that WebVoyager outperforms several baselines by leveraging both visual and textual signals. We also propose an automatic evaluation protocol by leveraging GPT-4V as the evaluator for online agents. Our work demonstrates the promise of using advanced LMM capabilities for building intelligent web agents. We hope WebVoyager provides a strong foundation for future research toward building more versatile and capable web assistants.
@@
Regarding the potential risks of our work, we believe that it requires a substantial amount of safety checks before deploying web agents like WebVoyagar into real-world applications, as the agent might unintentionally download malicious content from unauthorized websites, or input private/confidential information on public websites. Therefore it’s necessary to take extra caution when using and testing our agent.
@@ Appendix D Additional Trajectories
In our experiment, we present the complete trajectory of WebVoyager accessing the Apple website and addressing the task. In this section, we exhibit the trajectories for the remaining websites that complete web tasks successfully. We provide a screenshot for each step, accompanied by the action generated by WebVoyager. The specific navigation trajectories for each website are illustrated in Figures 9 to 22.
@@ Appendix E Additional Related Work
Vision-based Agents Concurrent to our work, a few related works also studied vision-based autonomous agents. VisualWebArena Koh et al. (2024) extends WebArena with additional websites and tasks that focus on visual reasoning to facilitate research on vision-based web agents. SeeClick Cheng et al. (2024) focused on finetuning a LMM to solely leverage screenshots as inputs to interact with websites. WebVLN Chen et al. (2023) introduced a web simulator that provides both HTML text and screenshots to finetune supervised vision-language models. GPT-4V Wonderland Yan et al. (2023) and AppAgent Zhang et al. (2023) instead focuses on building agents that can operate smart phone apps using the GPT-4V as the backbone. These works further underscore the promising prospects in this field.
@@ Appendix F Error Cases
In this section, we provide specific examples of the four types of errors mentioned in the Error Analysis section. Figure 23 illustrates a snippet of WebVoyager navigating on Google Flights and a Visual Grounding Issue appears. The task is to retrieve one-way flight information for January 22nd; however, it selects December 22nd on the Calendar and fails to make the necessary corrections. Although it attempts to modify the date in step 6, it ultimately fails to do so. Figure 24 illustrates a situation of WebVoyager navigating on Allrecipes, encountering the Navigation Stuck issue. The agent requires multiple downward scrolls to locate the correct ingredients. However, it experiences confusion during the process, uncertain whether to scroll up or down. Figure 25 depicts the Hallucination issue encountered by WebVoyager on the Coursera website. In the task, we query the number of quizzes in the "Artificial Intelligence for Healthcare" course. However, the agent only identifies the quiz in module 1 of the course, which is not the optimal answer and does not fulfill the task requirements. Figure 26 illustrates the issue of Prompt Misalignment encountered while browsing BBC News. WebVoyager fails to fulfill all the task requirements. Instead of completing the navigation, it provides partial answers and tells me how to find complete answers, which is not end-to-end.
@@
Figure 11: Screenshots of a complete trajectory of browsing ArXiv. Given the task: “For non-English submissions, do I need to provide a multi-language abstract, if need, answer the separator between the multiple abstracts.” The agent interacts with the ArXiv website and obtains the answer: “Line ("—–")”

### v4

@@ 1 Introduction
Figure 1: The overall workflow of WebVoyager. WebVoyager takes web tasks assigned by a human and automatically browses the web online. At each step, WebVoyager selects actions based on screenshots and text (the ‘type’ of the web element and its contents). Once the task is completed, the answers will be returned to the user. For example, for a user query: "Find the cost of a 2-year protection for PS4 on Amazon.", the agent interacts with Amazon online, locates the PS4, identifies the 2-year protection price, and returns "\$30.99" to the user.
@@
We conduct evaluations on a newly collected dataset, which is semi-automatically generated using a self-instruct (Wang et al., 2022) method, comprising 643 web tasks from 15 commonly accessed websites. We also evaluate WebVoyager on 90 web-related tasks of level 1 and level 2 from the GAIA (Mialon et al., 2023), and 50 interactive open-web tasks from SeeAct Zheng et al. (2024). We compare our WebVoyager with 1) GPT-4 (All Tools) [^2] GPT-4 (All Tools) is an integrated tool-based agent released by OpenAI in Oct. 2023. See https://chat.openai.com/, and 2) WebVoyager in a text-only setting which employs the textual accessibility tree proposed in WebArena (Zhou et al., 2023) to describe web pages. The results show that WebVoyager achieves a Task Success Rate of 59.1% on our new benchmark, significantly outperforming GPT-4 (All Tools) with a rate of 30.8% and the text-only setting with a rate of 40.1%, demonstrating the effectiveness of our method. Our research demonstrates the effectiveness of the WebVoyager method for web tasks, offering insights into the development of more intelligent and efficient web automation solutions.
@@ 2 Related Work
Along with these new benchmarks, numerous efforts have been made to build autonomous web agents. WebGPT (Nakano et al., 2021) constructs a text-based web browsing environment and fine-tunes GPT-3 as a web agent. WebAgent (Gur et al., 2023) pretrains a T5 model to extract HTML snippets and leverages Flan-U-PaLM (Chowdhery et al., 2023) to generate Python code to interact with the environment. Besides fine-tuning, another line of work tries to build web agents by prompting LLMs Yao et al. (2022b); Shinn et al. (2023); Ma et al. (2023). Multimodal web agents that integrate visual signals have also been explored, WebGUM (Furuta et al., 2023) combines T5 (Raffel et al., 2020) with a Vision Transformer (ViT) to navigate using both screenshots and HTML text. PIX2ACT (Shaw et al., 2023) instead solely relies on web screenshots as inputs to predict agent actions. Unlike previous works that only consider a single modality or simplified web environments, we build a multimodal agent that can complete tasks on real-world websites in this work. Concurrently with our work, SeeAct (Zheng et al., 2024) also leverages Large Multimodal Models (LMMs) for integrated visual understanding and actions on websites. However, the best SeeAct agent still relies on a finetuned cross-encoder model to select candidate elements for interaction. In contrast, WebVoyager does not require any additional modules.
@@ 3.3 Observation Space
Similar to how humans browse the web, our agent also takes the visual information from the web (screenshots) as the primary source of input. Using screenshots allows us to avoid the burden of processing HTML DOM tree or accessibility tree to portray the overall structure of webpages, which can lead to overly verbose texts and impact the decision-making process of the agent. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we overlay bounding boxes of the interactive elements on the websites to better guide the agent’s action prediction. Unlike Yang et al. (2023a), we do not need any object detection module (Zou et al., 2023). Instead, we utilize GPT-4V-ACT [^5] https://github.com/ddupont808/GPT-4V-Act, a Javascript tool to extracts the interactive elements based on web element types and then overlays bounding boxes with numerical labels on the respective regions of the elements. GPT-4V-Act is efficient since it is rule-based without incorporating any object detection model.
@@ 3.4 Action Space
We define the action space of our agent similar to how humans browse the web. To this end, we implement the most commonly used mouse and keyboard actions, sufficient for the agent to browse various web pages and locate the content required for the task. With the help of numerical labels in screenshots, we enable the agent to respond with a concise Action Format. This approach precisely locates the elements requiring interaction and executes the corresponding actions. The usage of actions is as follows (more details in Appendix C): 1) Click. This action involves clicking on an element within a webpage, typically a link or a button. 2) Input. This composite action involves selecting a text box, deleting any existing content within it, and then inputting new content. 3) Scroll. Scrolling is a common operation for browsing webpages, usually involving the vertical movement of the entire page. 4) Wait. Action execution requires time, and this action is often used to wait for web pages to load. 5) Back. This action is used to return to the previous page. 6) Jump to Search Engine. There are often situations where agents get stuck at a certain website without finding an answer. This action enables the agent to jump to a search engine and start anew. 7) Answer. Once all questions in the task are resolved, this action concludes the iteration and provides an answer in line with the task requirements.
@@ 4.2 Data Construction
To further confirm that the generated tasks in the dataset have low repetition, We use the all-mpnet-base-v2666https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model to calculate pairwise similarity for 643 ques. Out of a total of 206,403 pairs, only 49 pairs have a similarity greater than 0.8, and 140 pairs have a similarity between 0.7 and 0.8. All of these have been manually checked and found to be acceptable. 99.68% of pairs have a similarity of less than 0.6. This demonstrates the diversity of our tasks and the robustness of our approach.
@@ 5 Experiment
We employ GPT-4 Turbo with vision (gpt-4-vision-preview) as the backbone model of our agent, which showcases strong semantic and visual understanding capabilities equivalent to GPT-4V (OpenAI, 2023). Furthermore, we incorporate two additional backbone models, namely Claude 3 Opus (Anthropic, 2024) and GPT-4o (GPT-4 Omni, OpenAI, 2024), to augment the diversity of our research. We include the GPT-4 (All Tools) for baselines, which integrates vision, web browsing, code analysis, and various plugins in one model. Additionally, we consider a text-only baseline where the agent only receives the websites’ accessibility tree as input to prediction actions. We used a fixed browser window size of 1024 * 768 pixels for our environment, ensuring a consistent size for the screenshots in our observations. We set the temperature to 1 during generation and allow the agent to explore at most 15 steps.
@@ 5.1 Evaluation Methods
We adopt human evaluation as our main evaluation metric since most of the questions in our benchmark have open-ended answers. In particular, we provide the human evaluators with the complete trajectories of the agent’s interaction with the web (all screenshots and all actions), and ask them to provide a binary judgment of whether the agent successfully completed the task. For a subset of 300 tasks, we invite three annotators to judge each trajectory to understand the agreement among human annotators.
Even though human evaluations are accurate, they are often not scalable. Hence, we want to see if leveraging an LMM for automatic evaluation is feasible. To this end, we propose to use GPT-4V as an auto-evaluator that emulates the behavior of human evaluators to evaluate the navigation trajectories of WebVoyager. In particular, we provide the task, the responses from WebVoyager, and the last $k$ screenshots to the evaluator and ask it to judge whether the agent has completed the task, where $k$ is a hyper-parameter. The prompt of the GPT-4V evaluator is shown in Appendix B.
@@ 5.2 Result
We report the Agreement (the ratio of overlap) and Kappa ($\kappa$; Cohen 1960) between consolidated human labels777the Fleiss’s Kappa (Fleiss, 1971) of human annotators before any discussion is 0.7, which is substantial agreement. and GPT-4V’s judgments on the subset of 300 tasks in Table 2. Here, $k$ denotes the number of screenshots provided to GPT-4V, with “full” implying the full trajectory. GPT-4V’s agreement with human annotators gradually improves as it receives more information, and its final Kappa score also reaches 0.7, which is on par with the agreement among human annotators. The consistency between GPT-4V and humans suggests that GPT-4V is a promising automatic evaluator for multi-modal web agents. Accordingly, we report the automatic evaluation results based on GPT-4V in Table 1. The automatic evaluation results of three backbone models, GPT-4V, Claude 3 Opus, and GPT-4o, are relatively close, and their performance is significantly better than the Text-only setting (with GPT-4 as the backbone). However, there is a performance decline for both Claude and GPT-4o on Google Flights. Upon reviewing the trajectories, it is observed that GPT-4o falls into a cognitive bias, where it fails to correctly select the ‘one way’ option for one-way trip tasks, mistakenly assuming that only the departure date needs to be entered. On the other hand, Claude-3-Opus encounters difficulties in correctly interacting with web elements while inputting basic flight information. Modifying the system prompt for GPT-4o or Claude may potentially improve the performance.
WebVoyager Backbone Evaluator
GPT-4V Claude-3-Opus GPT-4o
GPT-4V 57.1 55.1 63.0
Claude-3-Opus 52.8 61.6 55.4
GPT-4o 55.5 54.9 64.1
Table 3: Overall Task Success Rate of WebVoyager using automatic evaluation. We employ GPT-4V, Claude-3-Opus, and GPT-4o as backbones for WebVoyager and run all tasks, followed by automatic evaluation using these three models.
Besides, we conduct the Claude-3-Opus based evaluation and the GPT-4o based evaluation. When provided with the full trajectory, the Claude-3-Opus achieves a kappa value of 0.6 with humans, indicating that it is less reliable than the GPT-4V. And the kappa value between GPT-4o and humans is 0.72, slightly higher than that of GPT-4V. Table 3 illustrates the Task Success Rate when using GPT-4V, Claude-3-Opus, and GPT-4o as backbones and auto evaluators. We observe that GPT-4o exhibits a more lenient attitude towards task performance results, while GPT-4V tends to be relatively strict. However, both models agree that Claude-3-Opus performs the worst. Claude-3-Opus, on the other hand, demonstrates a clear preference for its own results, believing that GPT-4V and GPT-4o are similar but considers itself to have the best performance. GPT-4o and GPT-4V also exhibit a certain bias towards their own results, with each considering itself to be superior to the other.
@@ 5.3 Discussions
Both text and vision are necessary for generalist web agents. As discussed earlier, WebVoyager struggles with text-heavy websites. On the other hand, we observe that the text-only agent demonstrates significantly poorer performance on websites with complex visual elements, such as Booking and Flights, which require interactions with calendars and other intricate components. In these scenarios, the textual input such as the accessibility tree becomes highly complex and verbose, making it far less intuitive than using screenshots. Hence it’s necessary to incorporate both modalities of inputs when building the general purpose agents.
@@
Why not use Open Source models. There are a few critical limitations of the existing open-sourced LMMs that prevent us from exploring other alternatives. Specifically, the web navigation task requires the agent to process fine-grained details from web page screenshots to make decisions, hence high-resolution is required to preserve the information from the web. However, most open-sourced LMMs such as LLaVA (Liu et al., 2024) reduce the image resolution to 224x224 or 336x336, which makes text with smaller fonts unrecognizable, hence they are unsuitable for web navigation tasks. Moreover, models like LLaVA have a max context length of 4096. In our case, the agent needs to handle trajectories as long as 15 steps, and it requires approximately 7000+ tokens, which does not fit in those models’ context size.
@@ 6 Conclusion
We introduce WebVoyager, an innovative web agent powered by large multimodal models (LMM) that can complete real-world web tasks end-to-end by interacting with websites. We have shown through evaluations that WebVoyager outperforms several baselines by leveraging both visual and textual signals. We also propose an automatic evaluation protocol by leveraging GPT-4V as the evaluator for online agents. Our work demonstrates the promise of using advanced LMM capabilities in building intelligent web agents. We hope WebVoyager provides a strong foundation for future research toward building more versatile and capable web assistants.
@@
Regarding the potential risks of our work, we believe that it requires a substantial amount of safety checks before deploying web agents like WebVoyagar into real-world applications, as the agent might unintentionally download malicious content from unauthorized websites, or input private/confidential information on public websites. Also the agent might send fake requests to website servers or generate fake user activities, which might be harmful to website owners. Therefore it’s necessary to take extra caution when using and testing our agent.
Ethics Statement
Our experiments have been designed to operate within strict ethical guidelines. Specifically, we restrict our web agent to perform only non-login tasks. This approach is in full compliance with the terms of service and user agreements of the websites our agent interacts with. Furthermore, we closely monitor the agent’s activities during its online evaluations. This monitoring is designed to identify and prevent any actions that could lead to potentially harmful consequences. By taking these precautions, we ensure that our research does not cross ethical boundaries or cause unintended harm.
Additionally, all task queries for evaluation undergo thorough manual inspection to ensure they are harmless and ethically sound. This manual inspection process is aimed at ensuring that every query is harmless and does not promote or propagate harmful content or actions. Our work aims to enhance user experience and accessibility while mitigating potential negative societal impacts. By proactively addressing ethics concerns, we dedicate to conducting research that benefits society while upholding high ethical standards.
@@ Appendix D Additional Trajectories
In Figure 4, we demonstrate how WebVoyager automatically navigates the Apple website and accomplishes the task. In this section, we exhibit the trajectories for the remaining websites that complete web tasks successfully. We provide a screenshot for each step, accompanied by the action generated by WebVoyager. The specific navigation trajectories for each website are illustrated in Figures 9 to 22. In addition, we also explore the performance of WebVoyager on websites in other languages, and we provide two example trajectories in Chinese and Spanish in Figure 23 and Figure 24.
@@ Appendix E Additional Related Work
Vision-based Agents Concurrent to our work, a few related works also studied vision-based autonomous agents. VisualWebArena Koh et al. (2024) extends WebArena with additional websites and tasks that focus on visual reasoning to facilitate research on vision-based web agents. SeeClick Cheng et al. (2024) focused on finetuning an LMM to solely leverage screenshots as inputs to interact with websites. WebVLN Chen et al. (2023) introduced a web simulator that provides both HTML text and screenshots to finetune supervised vision-language models. GPT-4V Wonderland Yan et al. (2023) and AppAgent Zhang et al. (2023) instead focus on building agents that can operate smartphone apps using the GPT-4V as the backbone. These works further underscore the promising prospects in this field.
@@ Appendix F Error Cases
In this section, we provide specific examples of the four types of errors mentioned in the Error Analysis section. Figure 25 illustrates a snippet of WebVoyager navigating on Google Flights and a Visual Grounding Issue appears. The task is to retrieve one-way flight information for January 22nd; however, it selects December 22nd on the Calendar and fails to make the necessary corrections. Although it attempts to modify the date in step 6, it ultimately fails to do so. Figure 26 illustrates a situation of WebVoyager navigating on Allrecipes, encountering the Navigation Stuck issue. The agent requires multiple downward scrolls to locate the correct ingredients. However, it experiences confusion during the process, and it is uncertain whether to scroll up or down. Figure 27 depicts the Hallucination issue encountered by WebVoyager on the Coursera website. In the task, we query the number of quizzes in the "Artificial Intelligence for Healthcare" course. However, the agent only identifies the quiz in module 1 of the course, which is not the optimal answer and does not fulfill the task requirements. Figure 28 illustrates the issue of Prompt Misalignment encountered while browsing BBC News. WebVoyager fails to fulfill all the task requirements. Instead of completing the navigation, it provides partial answers and tells me how to find complete answers, which is not end-to-end.
@@
Figure 11: Screenshots of a complete trajectory of browsing ArXiv. Given the task: “For non-English submissions, do I need to provide a multi-language abstract, if needed, answer the separator between the multiple abstracts.” The agent interacts with the ArXiv website and obtains the answer: “Line ("—–")”
@@
Figure 23: Screenshots of a complete trajectory of browsing Google Flights in Chinese. Given the task: “Find a flight from Hangzhou to Shenzhen on June 15, 2024, returning on June 20, departing at night, answer a flight that meets the requirements, and answer its specific departure time, airline and price.” The agent interacts with the Google Flights website and obtains the answer: “Departure time: 17:35, airline: Shenzhen Airlines, price: HK\$2,680”
Figure 24: Screenshots of a complete trajectory of browsing Cambridge Dictionary in Spanish. The description of task and answer are shown in Figure.